# Anisotropy of emergent large-scale dynamics in forced stratified shear flows

### v1

@@ 4 Discussion and perspective
Expecting dissipation to be dominated by the smallest scales in the flow, it might be surprising to see that dissipation (and mixing) statistics only converge for $L_{\mathrm{h,crit}}\gg\eta_{\textrm{K}}$ (with the Kolmogorov scale $\eta_{\textrm{K}}\sim\mathcal{O}\left(10^{-1}\right)$, see appendix B). Questioning this potential expectation, we plot the energy spectra associated with the dissipation rates of kinetic energy and buoyancy variance in figure 8 (c). Note that we show them in a pre-multiplied form (Krug et al., 2020) in order to highlight visually those wave numbers that cause most of the variance. We find the peaks in these spectra to be located at $\hat{k}\approx\left(3\dots 7\right)\times 10^{-1}$ or $\hat{\lambda}\approx 9\dots 18$, i.e., being again actually associated with the buoyancy interface depth or equivalently the turbulent mixing zone (due to the above-mentioned slaving). As we have shown in figure 8 (b) and explained in section 3.4, a convergence of this interface depth depends on resolving the streamwise spectral peak of the vertical buoyancy flux and requires $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$, explaining the convergence of the dissipation fields only in extended domains.
@@
Based on our spectral analysis from section 3.4, we have identified characteristic streamwise scales $\Lambda_{x}\approx 75\dots 115$ associated with the large-scale dynamics. We believe that this characteristic length scale can once again be interpreted in terms of the (sustained) emergent shear layer half-depth $\langle d\rangle_{t}\approx 8$ (which is measured in units of $d_{0}$). We remark that, although the imposed forcing is in principle designed to relax the flow back to the initial shear layer half-depth $d_{0}$ (and buoyancy interface half-depth $\delta_{0}$), the (above described) sustained turbulence results in or ensures this significantly deeper shear layer and buoyancy interface. It is well-known (see for example Scinocca & Ford (2000)) that the most unstable mode of KHI has a characteristic wavelength $\lambda_{\textrm{KHI}}$ of approximately fifteen times the shear layer (half-) depth. Here, at late times when $\langle d\rangle_{t}\approx 8$, this implies that $\lambda_{\textrm{KHI}}\sim 120$ which is very similar to the emergent streamwise length scales, particularly those associated with the streamwise velocity field as listed in table 2. However, this preferred instability scale is not able to roll up completely into coherent Kelvin-Helmholtz billows, and in particular, the turbulence certainly disrupts any possibility of subharmonic pairing or merging occuring, analogously to the disruption arguments put forward by Mashayek & Peltier (2013). Furthermore, our observation that spanwise scales are both significantly smaller and significantly harder to identify is consistent with the complete lack of billow coherence, as the knot/tube/defect structure is observed experimentally to require billows of at least $6003-5\lambda_{\textrm{KHI}}\sim 360-$ 600 spanwise extent, which is an order of magnitude larger than the observed spanwise scale $\Lambda_{y}\approx 50$.
In summary, our research leads to two main conclusions. Firstly, it adds to an increasing body of evidence that forced stratified shear flows ‘tune’ (in particular in the vertical direction) to a state with typical values of $\mathrm{Ri}_{\textrm{g}}\lesssim 0.2$. Such a state appears to allow both sustained turbulence and non-trivial buoyancy flux and attendant irreversible mixing in a fundamentally ‘weakly stratified’ regime. In this regime, buoyancy is slaved to the velocity field, and behaves, at least in some sense as a passive scalar, with relatively ‘inefficient’ mixing, in that bulk or average flux coefficients are bounded above by $0.2$, although occasional extreme events with substantially higher instantaneous local flux coefficients (see figure 5) can occur.
Secondly, it suggests the emergence of some imprint of the dominant (and inherently linear) shear-driven instability scale even in a turbulent flow. For numerical simulations to yield converged and reliable statistics therefore, domains should be scaled using the ‘tuned’ emergent vertical depth of the turbulent shear layer, which (during this tuning process) may be significantly larger than its initial value.
Clearly, these conclusions need to be tested for other choices of the key parameters. It would be very instructive to investigate the sensitivity of the ‘tuning’ process to the choice of the initial bulk Richardson number and Reynolds number. Moreover, an application to the ocean would clearly require the investigation of the sensitivity of the flow dynamics to the choice of $Pr$, as thermally stratified water has $Pr\sim\mathcal{O}\left(10\right)$. As we have seen, the emergent scales are both clearly related to the turbulent processes but also enormously larger than the smallest scales of that turbulence. A particular issue of interest is to consider larger initial $\mathrm{Re}_{0}$ that would be unstable right from the beginning of the simulation, avoiding the (potentially confusing) phenomenon of the shear layer effectively doubling in depth before the onset of the primary instability. If our conclusions prove to be robust and generic for other parameter choices, they have major implications for idealized computational studies, as they strongly suggest that to capture key characteristics of sheared turbulence in such ‘weak’ stratification, significantly extended computational domains are required.
There are also interesting implications for the geophysical application of such idealized studies of shear-driven stratified turbulence which completely ignore any effects of rotation. As our study suggests that streamwise scales can emerge that are $\mathcal{O}\left(50\right)$ times larger than the shear layer (half-depth) scale at initial onset of KHI, it is entirely plausible that such larger scales will at least be affected somewhat by rotational effects. As noted in the review of Taylor & Thompson (2023), ‘submesoscale motions’ with Rossby number $\mathrm{Ro}:=U/(fL)\sim O(1)$, where $f$ is the Coriolis parameter and $U$ and $L$ are characteristic velocity and length scales, are flows where ‘the Coriolis acceleration is important, but it does not constrain the motion’. In the world’s oceans, they ‘define submesoscales as dynamical features with horizontal scales between approximately 200 m and 20 km’. It is entirely plausible that geophysical flows such as estuarine outflows could have initial shear layer half-depths of the order of 4-10 m, and that KHI would onset immediately associated with that shear layer depth as observed for example by Holleman et al. (2016). In such flows, the emergent streamwise length scales may well experience non-trivial rotational effects. Therefore, it seems an interesting important question to investigate to what extent the emergent flow properties we have identified might be affected by larger-scale rotation.

### v2

@@ 4 Discussion
As presented in section 3 via a series of direct numerical simulations in domains of varying horizontal extent, the statistically stationary regime – which can be entered by stably stratified shear flows after an initial transient if subjected to a continuous forcing – can indeed be associated with a strongly hierarchical anisotropy of the finite-size large-scale dynamics. As shown in section 3.3, a convergence of vertical stirring- and mixing-related properties can be reached once the vertical extent of the mixing zone, as introduced in section 3.2, has converged. For the set of control parameters considered here, this has been the case for $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}=96$. However, expecting dissipation to be dominated by the smallest scales in the flow, it might be surprising to see that the dissipation (and mixing) statistics only converge for $L_{\mathrm{h,crit}}\gg\eta_{\textrm{K}}$ (with the Kolmogorov scale $\eta_{\textrm{K}}\sim\mathcal{O}\left(10^{-1}\right)$, see appendix A). Questioning this potential expectation, we plot the energy spectra associated with the dissipation rates of kinetic energy and buoyancy variance in figure 8 (c). Note that we show them in a pre-multiplied form (Krug et al., 2020) in order to highlight visually those wave numbers that cause most of the variance. We find the peaks in these spectra to be located at $3\times 10^{-1}\lesssim\hat{k}\lesssim 7\times 10^{-1}$ or $9\lesssim\hat{\lambda}\lesssim 18$, i.e., being again actually associated with the depth of the turbulent mixing zone. As we have shown in figure 8 (b) and explained in section 3.4, a convergence of this depth depends on resolving the streamwise spectral peak of the vertical buoyancy flux and requires $L_{\mathrm{h}}\gtrsim L_{\mathrm{h,crit}}$. Hence, the convergence of statistical aspects of dissipation goes hand in hand with the macroscopic or structural convergence of the turbulent mixing zone.
This convergence of the depth of the turbulent mixing zone to a finite characteristic value $\Lambda_{z}=2\langle d\rangle_{t}\gg 2$ (remember that we measure lengths in units of $d_{0}$) for sufficiently extended domains, despite the ongoing forcing and mixing, is a central result of section 3.2. Before this convergence was reached, we observed a growth of its half-depth with an increasing extent of the domain. This growth can be explained by an increased number of degrees of freedom of the underlying dynamical system, leading to a higher potential complexity, and thus more vigorous turbulence that causes both a stronger macroscopic stirring as well as a stronger microscopic mixing. However, this trend must eventually be limited at some point as such stirring increases the strength of the stratification.
Simply, the deepening of the mixing zone cannot continue without limit since turbulence cannot be sustained when the stratification becomes too strong. This strength of stratification can be quantified by either the bulk Richardson number $\mathrm{Ri}$ or appropriate averages of the gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$, see again figure 4 (c). The particular observation that the (mixing zone) average of $\mathrm{Ri}_{\textrm{g}}$ is approximately $0.15$ is not only supported by the Miles-Howard criterion but also highly reminiscent of the results reported by Portwood et al. (2019). They used a control strategy (effectively through modulating gravity) to identify equilibrium turbulent states in linearly stratified flows driven by constant vertical shear. Although the flow considered here is different, the key dynamics appears to be similar: the forced flow adjusts until the stratification is as strong as possible to still allow for vigorous turbulence which is able to stir and mix the buoyancy field.
Hence, we hypothesise that forced stratified shear flows ‘tune’ themselves towards equilibrium states associated with typical values of $\mathrm{Ri}_{\textrm{g}}\lesssim 0.2$, eventually governing the resulting depth of the turbulent mixing zone. This ‘tuning’ is highly reminiscent of the ‘kind of equilibrium’ proposed by Turner (1973), and indeed has points of connection with the more recently-proposed arguments of ‘self-organised criticality’ in stratified shear flows (Salehipour et al., 2018; Smyth et al., 2019). This tuning was also observed in wall-bounded stratified plane Couette flow by Zhou et al. (2017), where analogously, an apparently maximum possible midplane gradient Richardson number for sustained turbulence $\mathrm{Ri}_{\textrm{g}}\simeq 0.2$ was identified. Zhou et al. (2017) characterised the closeness of this number to the marginal Miles-Howard criterion of $1/4$ as being potentially ‘fortuitous’, and argued that it may be more appropriate to interpret this value in terms of emergent turbulent balances, consistently with the deep physical insights of Turner (1973).
@@
To complement this identification of a characteristic vertical scale associated with the large-scale dynamics, we have also identified characteristic streamwise scales $75\lesssim\Lambda_{x}\lesssim 115$ associated with the large-scale dynamics. We believe that this characteristic length scale can once again be interpreted in terms of the (sustained) emergent shear layer half-depth $\langle d\rangle_{t}\approx 8$. We remark that, although the imposed forcing is in principle designed to relax the flow back to the initial shear layer half-depth $d_{0}$ (and buoyancy interface half-depth $\delta_{0}$), the (above described) sustained turbulence ensures this long-time significantly deeper shear layer and buoyancy interface. It is well-known (see for example Scinocca & Ford (2000)) that the most unstable mode of KHI has a characteristic wavelength $\lambda_{\textrm{KHI}}$ of approximately fifteen times the shear layer half-depth. Given $\langle d\rangle_{t}\approx 8$ at late times, this implies $\lambda_{\textrm{KHI}}\approx 120$ which is very similar to the emergent streamwise length scales, particularly those associated with the streamwise velocity field as listed in table 2.
This observation encourages us to conduct a classical (unforced) linear stability analysis – following the procedures outlined in (Smyth et al., 2011; Smith et al., 2021) – of the emergent background profiles $\langle u_{x}\rangle_{A,t}$ and $\langle b\rangle_{A,t}$ together with the mean associated turbulent diffusivities of momentum and buoyancy as built on the approach of Kaminski & Smyth (2019). The required decomposition of the background flow from the turbulent fluctuations is discussed in more detail below. Indeed, it turns out that the mode of fastest (real) growth rate is associated with a wavelength $109\lesssim\lambda_{\textrm{background KHI}}\lesssim 121$. This highlights that even a sustained turbulent interface can be susceptible to a Kelvin-Helmholtz instability, with crucial implications for the emergent large-scale dynamics.
However, this preferred instability scale is not able to roll up completely into coherent Kelvin-Helmholtz billows, and in particular, the turbulence certainly disrupts any possibility of subharmonic pairing or merging occuring, perhaps analogously to the disruption arguments put forward by Mashayek & Peltier (2013). Furthermore, our observation that spanwise scales are both significantly smaller and significantly harder to identify is consistent with the complete lack of billow coherence, as the knot/tube/defect structure is observed experimentally to require billows of at least $3-5\lambda_{\textrm{KHI}}\sim 360-600$ spanwise extent, which is an order of magnitude larger than the observed spanwise scale $\Lambda_{y}\approx 50$. We remark that, both due to their temporal stationarity and absolute size, the here extracted large-scale dynamics appears to differ significantly and qualitatively from the inherently transient structures that monotonically ‘increase with time approximately from $5d_{0}$ to $9d_{0}$ in an unforced configuration as described in (Watanabe & Nagata, 2021).
Figure 9: Additional characteristics of the emergent dynamics. The mixing zone is characterised by: (a) heightened values of $\mathrm{Re}_{\textrm{b}}\gtrsim 35$ and $\mathrm{Fr}$; (b) values of the bulk mixing coefficient smaller than the canonical (bounding) value of $\bar{\Gamma}_{\chi}=0.2$ (marked by the vertical line). (c) A comparison of present dynamically manifesting and global time scales highlights that the vast scale separation in space is complemented by another one in time as well. Here, the flow has $L_{\mathrm{h}}=128$, similarly to figure 3.
The sustained large-scale dynamics is associated with a – relative to its initial value or profile – significantly increased gradient Richardson number $\mathrm{Ri}_{\textrm{g}}$ as shown in figure 3 (b). However, this single measure is (of course) insufficient to characterise stratified shear flows. Two other commonly used parameters quantifying the relative strengths of turbulence, stratification and viscosity are the buoyancy Reynolds number
$\mathrm{Re}_{\textrm{b}}\left(z\right):=\mathrm{Re}_{0}\frac{\langle\varepsilon_{u}\rangle_{A,t}}{\langle N^{2}\rangle_{A,t}}\equiv\left(\frac{t_{\textrm{N}}}{t_{\textrm{K}}}\right)^{2}$ (24)
and the (turbulent) Froude number
$\mathrm{Fr}\left(z\right):=\frac{\langle\varepsilon_{u}\rangle_{A,t}}{\langle E_{\textrm{kin}}\rangle_{A,t}\langle N\rangle_{A,t}}\equiv\frac{t_{\textrm{N}}}{t_{\textrm{turb}}}$ (25)
where $E_{\textrm{kin}}:=\left(u_{x}^{2}+u_{y}^{2}+u_{z}^{2}\right)/2$ is the (pointwise and total) kinetic energy (density). Figure 9 (a) shows the vertical profiles of these two parameters $\mathrm{Re}_{\textrm{b}}$ and $\mathrm{Fr}$ for the same flow as shown for $\mathrm{Ri}_{\textrm{g}}$ in figure 3 (b).
Interestingly, this figure suggests an (effectively equivalent) alternative interpretation of the mixing zone as the region where the buoyancy Reynolds number has a heightened value of $\mathrm{Re}_{\textrm{b}}\gtrsim 35$. We remark that $\mathrm{Re}_{\textrm{b}}$ can also be written as a ratio of length scales $\mathrm{Re}_{\textrm{b}}\equiv\left(\eta_{\textrm{O}}/\eta_{\textrm{K}}\right)^{4/3}$ where $\eta_{\textrm{K}}$ is the Kolmogorov microscale and $\eta_{\textrm{O}}:=\sqrt{\langle\varepsilon_{u}\rangle_{A,t}/\langle N^{3}\rangle_{A,t}}$ is the Ozmidov length scale. Here $\eta_{\textrm{K}}$ is assumed to vary with height (as introduced and plotted in appendix A) and thus calculated based on $\langle\varepsilon_{u}\rangle_{A,t}$. Since $\eta_{\textrm{O}}$ can be interpreted as the characteristic scale of the largest eddy not significantly affected by the stratification whereas $\eta_{\textrm{K}}$ is the scale at which viscous dissipation completely damps motions, such larger values of $\mathrm{Re}_{\textrm{b}}$ give at least some opportunity of a range of turbulent scales neither strongly affected by stratification nor by viscosity and thus allow (potentially) for vigorous turbulence and mixing. Dating back to the seminal work of Gibson (see for example Gibson (1987)), there is much evidence that such values of $\mathrm{Re}_{\textrm{b}}$ are associated with a recognisably turbulent stratified mixing dynamics, see for example Smyth & Moum (2000) for a further discussion. Indeed, Shih et al. (2005) argued that $7\lesssim\mathrm{Re}\lesssim 100$ defined a transitional mixing regime with essentially constant bulk flux coefficient close to the canonical value of 0.2.
We also see that the mixing zone is associated with heightened values of $\mathrm{Fr}$, perhaps unsurprising after consideration of figure 3 (b) since there is a commonly assumed correlation between $\mathrm{Ri}_{\textrm{g}}\propto\mathrm{Fr}^{-2}$. However, caution should be exercised when considering particular numerical values of $\mathrm{Fr}$ as $\mathrm{Ri}_{\textrm{g}}\lesssim 0.25$ in the mixing zone, and so it is appropriate to think of the (shear-driven) mixing to be relatively weakly stratified and ‘dominated’ by the shear (Mater & Venayagamoorthy, 2014).
Furthermore, as we find an average of $\eta_{\textrm{O}}\approx 6$ across the turbulent mixing zone, we recover the (length scale) hierarchy $\eta_{\textrm{K}}\ll\eta_{\textrm{O}}\ll\Lambda_{z}\ll\Lambda_{y}\ll\Lambda_{x}$, entirely consistently with the observation that there are no large-scale overturnings of the entire mixing zone. Although this observed hierarchy of scales in this self-organising emergent flow is consistent with that observed in the strongly stratified regime (sometimes also called the layered anisotropic stratified turbulence or ‘LAST’ regime) considered by Brethouwer et al. (2007), the dominant cause for anisotropy in this flow is the continually imposed shear forcing which leads (at least within the mixing zone) to relatively weakly stratified shear-driven (and shear-dominated) turbulence.
Quantitative evidence of this relatively weakly stratified shear-driven turbulence can be obtained from consideration of mixing quantities. In section 3.3 we introduced the (point-wise determined) local flux coefficient $\Gamma_{\chi}$ to highlight that even statistical aspects of the flow are affected by the large-scale dynamics. However, it is more common to describe this ratio of dissipation rates via average values of $\varepsilon_{u}$ and $\chi$. Hence, we additionally define a bulk mixing coefficient
$\bar{\Gamma}_{\chi}:=\frac{\langle\chi\rangle_{A,t}}{\langle\varepsilon_{u}\rangle_{A,t}}$ (26)
and plot its vertical profile in figure 9 (b). Note the relation of this panel to figure 3 (d). Using the approach described in section 3.3, we find $\bar{\Gamma}_{\chi}\left(z=0\right)\simeq 0.08$ at the midplane and $\langle\bar{\Gamma}_{\chi}\rangle_{-\langle d\rangle_{t}\leq z\leq\langle d\rangle_{t}}\simeq 0.13$ within the turbulent mixing zone, essentially because $\varepsilon_{u}$ is more localised at the midplane than $\chi$ is. These values are somewhat smaller than the canonical value of $0.2$ proposed by Osborn (1980), though it is important to remember that this canonical value is a maximum rather than the more commonly assumed constant value. Within Osborn’s arguments this maximum value of $\Gamma\leq 0.2$ is associated with typical empirically observed maximum values of the ‘flux’ Richardson number $\mathrm{Ri}_{\textrm{f}}\simeq\bar{\Gamma}_{\chi}/(1+\bar{\Gamma}_{\chi})$, implying $\mathrm{Ri}_{\textrm{f}}\lesssim 0.17$ within our mixing zone. As demonstrated for example by Zhou et al. (2017) – see also more detailed and pedagogical discussions in the review of Caulfield (2021) –, $\mathrm{Ri}_{\textrm{f}}\simeq\mathrm{Ri}_{\textrm{g}}$ for weakly stratified shear-driven mixing and so this maximum value can be thought of as an appropriate maximum value for the ‘gradient’ Richardson number as well. As shown in figure 4 (c), $\mathrm{Ri}_{\textrm{g}}\simeq 0.15$ when averaged across the mixing zone and so it would seem superficially reasonable that $\bar{\Gamma}_{\chi}$ should be larger, since if $\mathrm{Ri}_{\textrm{f}}\simeq\mathrm{Ri}_{\textrm{g}}$ then $\bar{\Gamma}_{\chi}\simeq\mathrm{Ri}_{\textrm{g}}/(1-\mathrm{Ri}_{\textrm{g}})\simeq 0.18$.
However, there are at least two plausible reasons for the discrepancy. First, and perhaps most importantly, central to Osborn’s model is the assumption that mixing (however quantified) is linearly proportional to the dissipation of turbulent kinetic energy. In our flows, it is clear that although both $\varepsilon_{u}$ and $\chi$ are elevated in the mixing zone, the assumption that they are linearly related is clearly neither qualitatively nor quantitatively correct. The second reason is somewhat more subtle and related to the specific properties of the flow forcing being used by us here. As $\chi$ actually involves diffusion explicitly, it seems most natural to define a flux coefficient in terms of $\chi$ (and potentially its averages) rather than in terms of the buoyancy flux B-B which is not even sign-definite, although Osborn (1980) did define the flux coefficient using B-B in the numerator. The two quantities B-B and $\chi$ balance each other in an appropriately constructed buoyancy variance equation when there is no forcing as discussed, for example, in Caulfield (2021), making the choice of $\chi$ or B-B in the numerator of $\Gamma$ moot. However, this balance does not occur in the presence of forcing as is apparent from the data shown in figures 3 (c, d). For the present flow, we find $\langle B\rangle_{A,t}/\langle\chi\rangle_{A,t}\approx 3$ at midplane or $\approx 1.5$ when averaged across the turbulent mixing zone. This mismatch is entirely due to the particular form of the forcing which we are using. The forcing of both the velocity and buoyancy fields to relax back to their original profiles inevitably perturbs in nontrivial ways both the kinetic energy and the potential energy, as well as (crucially) the direct exchange between these two reservoirs, and hence the buoyancy flux. This (admittedly somewhat artificial) observation underlines that, for our forced configuration, $\chi$ – which naturally continues to describe the irreversible mixing processes – is the appropriate reference for the definition of a flux coefficient $\Gamma$, although this choice may contribute to the numerical discrepancy between our observed $\bar{\Gamma}_{\chi}$ and expected value within the theoretical framework of Osborn (1980).
Throughout this work, when considering the flow dynamics, we have always used the total solution fields. (Furthermore, we have defined buoyancy in terms of density deviations from the constant reference density rather than relative to the statically stable density field.) An alternative perspective is to decompose the fields into some background (or appropriately averaged) flow and the residual turbulent fluctuations, i.e., $\Phi=\bar{\Phi}+\Phi^{{}^{\prime}}$ where the overbar denotes an average over which the fluctuations vanish. For our configuration, the appropriate decomposition is
$u_{x}$ $=\bar{u}_{x}$ $+$ $u_{x}^{{}^{\prime}}$ with $\thickspace\thickspace\bar{u}_{x}\left(z\right)$ $:=\langle u_{x}\rangle_{A,t},$ (27)
$u_{y}$ $=$ $u_{y}^{{}^{\prime}},$ (28)
$u_{z}$ $=$ $u_{z}^{{}^{\prime}},$ (29)
$b$ $=\bar{b}$ $+$ $b^{{}^{\prime}}$ with $\thickspace\thickspace\bar{b}\left(z\right)$ $:=\langle b\rangle_{A,t}$ (30)
where the time-independent background profiles emerge dynamically due to the continuous forcing defined in equations (6) and (7). Considering the dissipation rates as defined in equations (16) and (17), this leads to
@@
[…]
with the strain rate tensor $\bm{S}=\bar{\bm{S}}+\bm{S}^{{}^{\prime}}$. By definition, the mixed contributions vanish on average, i.e. $\langle\varepsilon_{u}^{\times}\rangle_{A,t}=0$ and $\langle\varepsilon_{b}^{\times}\rangle_{A,t}=0$. In contrast, the vast majority of the dissipation is associated with the turbulent fluctuations. We find that $\langle\varepsilon_{u}^{{}^{\prime}}\rangle_{A,t}/\langle\varepsilon_{u}\rangle_{A,t}\approx\langle\varepsilon_{b}^{{}^{\prime}}\rangle_{A,t}/\langle\varepsilon_{b}\rangle_{A,t}\gtrsim 0.8$ at all heights and the minimum happening at the midplane, perhaps unsurprisingly as that is the location of the largest vertical gradient in the background fields. Hence, these observations give us confidence that our calculations based on the total fields (incorporated, e.g., in $\mathrm{Ri}_{\textrm{g}}$, $\mathrm{Re}_{\textrm{b}}$, and $\mathrm{Fr}$) enable equivalent insights for the turbulent fields.
Finally, in order to highlight the (sub-)dominance of various mechanisms in our forced stratified shear flow, we compare various time scales in figure 9 (c). Note that, based on our non-dimensionalisation from equation (1), all our non-dimensional time scales are related via $\tau_{\Phi}=t_{\Phi}\tau_{\textrm{adv}}$ to their dimensional counterpart and, thus, scaled relative to the dimensional advective time scale $\tau_{\textrm{adv}}$ of the background shear flow. We consider the Kolmogorov, shear, buoyancy frequency, and turbulence decay time scales – defined by
$t_{\textrm{K}}:=\frac{1}{\sqrt{\mathrm{Re}_{0}\langle\varepsilon_{u}\rangle_{A,t}}},\quad t_{\textrm{S}}:=\frac{1}{\langle S\rangle_{A,t}},\quad t_{\textrm{N}}:=\frac{1}{\langle N\rangle_{A,t}},\thickspace\textrm{and}\quad t_{\textrm{turb}}:=\frac{\langle E_{\textrm{kin}}\rangle_{A,t}}{\langle\varepsilon_{u}\rangle_{A,t}},$ (33)
respectively – as those time scales that are associated with the emergent flow. The emergent non-dimensional (and vertically varying) parameters $\mathrm{Ri}_{\textrm{g}}$, $\mathrm{Re}_{\textrm{b}}$, and $\mathrm{Fr}$ can then be interpreted as ratios of these timescales as shown on the right of equations (13), (24), and (25). Conversely, the advective, viscous diffusion, (bulk) buoyancy, and relaxation time scales – defined by
$t_{\textrm{adv}}:=1,\quad t_{\nu}:=\mathrm{Re}_{0},\quad t_{\textrm{b}}:=\frac{1}{\mathrm{Ri}_{0}},\thickspace\textrm{and}\quad t_{\textrm{r}},$ (34)
respectively – represent ‘global’ time scales which are determined by the externally imposed control parameters. In general, the Batchelor and the buoyancy diffusion time scale, $t_{\textrm{B}}:=t_{\textrm{K}}/\sqrt{\mathrm{Pr}}$ and $t_{\kappa}:=\mathrm{Re}_{0}\mathrm{Pr}$, would be two further important time scales. However, since $\mathrm{Pr}\equiv t_{\kappa}/t_{\nu}=\left(t_{\textrm{K}}/t_{\textrm{B}}\right)^{2}=1$, $t_{\textrm{B}}=t_{\textrm{K}}$ as well as $t_{\kappa}=t_{\nu}$ here. Figure 9 (c) highlights a hierarchy $t_{\textrm{adv}}<t_{\textrm{K}}<t_{\textrm{S}}<t_{\textrm{N}}<t_{\nu}<t_{\textrm{b}}<t_{\textrm{r}}$ across the entire turbulent mixing zone while $t_{\textrm{turb}}\approx t_{\nu}$ at midplane but $t_{\textrm{r}}\ll t_{\textrm{turb}}$ with increasing distance. As processes associated with shorter time scales tend to be stronger, this underlines that any effects introduced by the volumetric forcing can be considered sub-dominant. Moreover, finding $t_{\textrm{S}}<\left\{t_{\textrm{N}},t_{\textrm{turb}}\right\}$ underlines yet again that our forced flow lies in a weakly stratified shear-dominated regime (Mater & Venayagamoorthy, 2014; Caulfield, 2021).
5 Conclusions
Motivated by their relevance to many geophysical flows, we have conducted direct numerical simulations of forced stratified shear flows to improve our fundamental understanding of their dynamics. Although our flow configuration is prone to a primary Kelvin-Helmholtz instability, it additionally and importantly includes a continuous forcing which tends to restore the initial profiles and ensures that the ensuing turbulence can be sustained for arbitrarily long times. Since it is this continuous forcing which connects to the motivating geophysical flows, we largely disregard the initial transient but rather focus on the regime of statistically stationary dynamics during later times.
One of our objectives was to determine whether or not the resulting shear layer depth $d$ converges to a finite value – which is independent of the extent of the domain and thus characteristic of the emergent self-organisation of the flow – despite the ongoing forcing and mixing. To this end, in section 3, we have simulated this regime in numerical domains of varying horizontal extent in the range $16\leq L_{\mathrm{h}}\leq 512$ (measured in units of the initial shear layer half-depth $d_{0}$). We found that the resulting total shear layer depth indeed converges to a finite value $\Lambda_{z}=2\langle d\rangle_{t}\approx 16$ once $L_{\mathrm{h}}\gtrsim 96$. Moreover, we have also identified characteristic streamwise and spanwise length scales – $75\lesssim\Lambda_{x}\lesssim 115$ and $\Lambda_{y}\approx 50$, respectively – associated with the emergent large-scale dynamics. As discussed in more detail in section 4, the emergent dynamics of this sustained turbulent flow can be associated with a hierarchy of length and time scales – $\eta_{\textrm{K}}\ll\eta_{\textrm{O}}\ll\Lambda_{z}\ll\Lambda_{y}\ll\Lambda_{x}$ and $t_{\textrm{adv}}<t_{\textrm{K}}<t_{\textrm{S}}<t_{\textrm{N}}<t_{\nu}\lesssim t_{\textrm{turb}}<t_{\textrm{b}}<t_{\textrm{r}}$, respectively –, underlining that our self-organising emergent flow demonstrates strongly anisotropic behaviour.
In summary, our research leads to two main conclusions. Firstly, as forced stratified shear flows result in a significantly enlarged yet finite shear layer depth, it adds to an increasing body of evidence that such flows ‘tune’ (in particular in the vertical direction) to a quasi-equilibirum state with typical values of $\mathrm{Ri}_{\textrm{g}}\lesssim 0.2$. Such a state appears to allow both sustained turbulence and non-trivial buoyancy flux and attendant irreversible mixing in a fundamentally ‘weakly stratified’ (and hence, due to the forcing, shear-dominated) regime. Secondly, as the most unstable mode of the emergent background flow has a length scale that is very similar to the emergent length scale of the characteristic large-scale dynamics of the streamwise velocity field, we conjecture that even such strongly turbulent flows as the flows consider here may retain an imprint of the (inherently linear) shear-driven primary instability scale. From a more practical perspective, these two conclusions imply that domains should be scaled using the finite ‘tuned’ emergent vertical depth of the turbulent shear layer, which (during this tuning process) may be significantly larger than its initial value in order to yield converged and reliable statistics. If the large-scale pattern formation is of interest, the streamwise and spanwise extents may need to be even larger, requiring highly extended domains.
Clearly, these conclusions need to be tested for other choices of the key parameters. It would be very instructive to investigate the sensitivity of the ‘tuning’ process to the choice of the initial bulk Richardson number and Reynolds number. Moreover, an application to the ocean would clearly require the investigation of the sensitivity of the flow dynamics to the choice of $\mathrm{Pr}$, as thermally stratified water has $\mathrm{Pr}\sim\mathcal{O}\left(10\right)$. As we have seen, the emergent scales are both clearly related to the turbulent processes but also enormously larger than the smallest scales of that turbulence. A particular issue of interest is to consider larger initial $\mathrm{Re}_{0}$ that would be unstable right from the beginning of the simulation, avoiding the (potentially confusing) phenomenon of the shear layer effectively doubling in depth due to molecular diffusion before the onset of the primary instability. If our conclusions prove to be robust and generic for other parameter choices, they have major implications for idealized computational studies, as they strongly suggest that to capture key characteristics of sheared turbulence in such relatively ‘weak’ stratification, significantly extended computational domains are required. Furthermore, we have focussed exclusively on the end-member case of a continually forced shear layer, in contrast to the more widely-studied other end-member case of an unforced shear layer undergoing a ‘run-down’ mixing event. Real flows will in general exhibit a transient forcing that lies somewhere between these two end members, and it is an interesting and open question as to how strongly dependent the flow dynamics would be on a finite time scale of forcing.
There are also interesting implications for the geophysical application of such idealized studies of shear-driven stratified turbulence which conventionally completely ignore any effects of rotation. As our study suggests that streamwise scales can emerge that are $\mathcal{O}\left(50\right)$ times larger than the half-depth of the shear layer during the initial onset of KHI, it is entirely plausible that such larger scales will at least be affected somewhat by rotational effects. As noted in the review of Taylor & Thompson (2023), ‘submesoscale motions’ with Rossby number $\mathrm{Ro}:=U/(fL)\sim O(1)$, where $f$ is the Coriolis parameter and $U$ and $L$ are characteristic velocity and length scales, are flows where ‘the Coriolis acceleration is important, but it does not constrain the motion’. In the world’s oceans, they ‘define submesoscales as dynamical features with horizontal scales between approximately 200 m and 20 km’. It is entirely plausible that geophysical flows such as estuarine outflows could have initial shear layer half-depths of the order of 4-10 m, and that KHI would onset immediately associated with that shear layer half-depth as observed for example by Holleman et al. (2016). In such flows, the emergent streamwise length scales may well experience non-trivial rotational effects. Therefore, it seems an interesting important question to investigate to what extent the emergent flow properties we have identified might be affected by larger-scale rotation.
[…]